6. Bovous Stores, Inc., sells several products. Information of average revenue and costs is as follows: Selling price per unit $20.00 Variable costs per unit: Direct material $4.00 Direct manufacturing labor $1.60 Manufacturing overhead $0.40 Selling costs $2.00 Annual fixed costs $96,000 The revenues that the company must earn annually to make a profit of $144,000 are: A) $378,000 B) $425,000 C) $400,000 D) $450,000
Answer #1

Calculation of revenue required to make profit:

Profit = $144000

Fixed cost= $96000

Variable cost per unit= 4+1.6+0.4+2= 8

Selling price= 20

Contribution per unit= 20-8= 12

Contribution margin ratio= 12/20= 0.60

Revenue= (fixed cost+ profit)/contribution margin ratio

Revenue= (96000+144000)/0.60= 240000/0.60= 400000

So correct answer is C) $400000